# Bubble Nucleation from Boson Star Collapse

### v1

@@ Abstract
We present a new classical mechanism for nucleation of bubbles of true vacuum. The mechanism arises when dense boson stars form in the false vacuum. As the boson stars collapse due to attractive self-interactions, the field inside the stars core is enhanced beyond the potential barrier. Subsequently the stars explode as true vacuum bubbles, and induce a cosmological phase transition. The mechanism raises the possibility that a vacuum that is stable against quantum tunneling can be vulnerable to “astrophysical” processes.
@@ 1 Introduction
Recent advances in gravitational wave astronomy Barausse:2020rsu; NANOGrav:2023hvm have sparked interest in the physics of cosmological first-order phase transitions Coleman:1977py; Linde:1980tt, potentially leading to stochastic gravitational wave signals that are observable in current and future experiments. These experiments can reach frequencies up to a few kHz, which correspond to transition scales much higher than the physics scales reachable by collider experiments. [^1] A present-day frequency of kHz is tied to a transition scale of $\sim 10^{10}\,\mathrm{GeV}$ in conventional scenarios. This relation however is modified in the mechanism we will present. The signal details depend strongly on the physics behind the phase transition, which usually proceeds via quantum tunnelling Coleman:1977py or thermal fluctuations Linde:1980tt; LINDE1980289 creating supercritical bubbles which later expand and collide. In this paper, we would like to present a new classical mechanism for the creation of supercritical bubbles from the collapse of a boson star.
Boson stars Kaup:1968zz; Ruffini:1969qy are clumps of Bose-Einstein condensate that are held in equilibrium by gravity or attractive self-interactions. They are known to arise in various early universe scenarios Colpi:1986ye; Liddle:1992fmk; Kolb:1993zz; Khlebnikov:1999pt; Sikivie:2009qn; Chavanis:2011zi; Chavanis:2011zm; Erken:2011vv; Davidson:2013aba; Davidson:2014hfa; Guth:2014hsa; Eby:2015hsq; Levkov:2016rkk; Schiappacasse:2017ham; Levkov:2018kau; Eggemeier:2019jsu; Chen:2021oot; Chan:2022bkz; Dmitriev:2023ipv; Gorghetto:2024vnp; Ralegankar:2024zjd, and their physics has been extensively discussed especially in the context of axions (see e.g. Marsh:2015xka; Zhang:2018slz for reviews). One of the important consequences of boson stars is that their evolutionary track can end with a collapse, which drives the energy density in the core of the star to very large values. This is indeed the case for axions, whose attractive self-interaction renders axion stars unstable against collapse; the axion star collapse has been shown Levkov:2016rkk to be accompanied by emission of relativistic axions from the high-density core.
@@
Generally speaking, the process we investigate can be viewed as a true vacuum bubble creation driven by field–interaction dynamics. In this context, we would like to mention Refs. Easther:2009ft; Giblin:2010bd, where a different type of field dynamics also leads to classical bubble production. In those works, large field excursions are generated by bubble collisions, which in turn result in the classical creation of new bubbles in minima lying below those of the parent bubbles. However, unlike the mechanism discussed in the present paper, the effect found in Easther:2009ft; Giblin:2010bd operates only in models featuring a sequence of at least two false vacua arranged in a hierarchical structure. Finally, let us also mention other studies that are somewhat similar in spirit, where phase transitions are triggered by various types of nucleation sites Hosotani:1982ii; Preskill:1992ck, such as neutron stars Balkin:2021wea; Balkin:2021zfd, domain walls Blasi:2022woz, cosmic strings Yajnik:1986tg; Blasi:2024mtc, and monopoles Steinhardt:1981ec; Kumar:2010mv; Agrawal:2022hnf.
The paper is structured as follows: in Section 2 we briefly review dense boson stars, focusing on their collapse according to the self-similar solution found by Levkov:2016rkk. In Section 3, we present the main result of this work and show for an explicit potential, how the boson star collapse can trigger the phase transition. In Section 4, we comment on phenomenological applications of this mechanism. Finally in Section 5 we discuss our findings and provide with our concluding remarks. We also give a general discussion about the critical radius of a true vacuum bubble in Appendix A.
@@ 2 Boson stars: a review
Let us start by reviewing the basic properties of boson stars, which are chunks of Bose-Einstein condensate bound by gravity or attractive self-interactions. Some of the discussions here follow those of Guth:2014hsa; Levkov:2016rkk; Schiappacasse:2017ham.
@@ 2.1 Stable and unstable branches
where $r=\lvert\bm{x}\rvert$. Then the integrals in (7) can be performed to yield [^3] For exact star solutions of the GPP equations, the terms in the Hamiltonian (9) can be corrected by order-unity factors, whose values also depend on the detailed definitions of the star radius and mass Ruffini:1969qy; Membrado:1989ke; Visinelli:2017ooc. However the approximation (8) is good enough for order-of-magnitude estimates.
@@ 2.2 Boson star collapse
We now focus on star configurations in the vicinity of the unstable branch, where gravity can be ignored. Then the GPP equations (6) reduce to a nonlinear Schrödinger-type equation, which has been well studied in condensed matter systems Zakharov-Kuznetsov,
@@
As noted in Zakharov-Kuznetsov; Levkov:2016rkk, this equation is invariant under the scaling,
@@
where $A$ and $B$ are complex constants. The second term should vanish for the Hamiltonian to be finite. Numerically solving (15) under these boundary conditions fix the parameters as $w\simeq 0.54$ and $\lvert A\rvert\simeq 2.85$ Levkov:2016rkk. Moreover, the absolute value of $\chi$ thus obtained is well approximated by the following piece-wise function,
@@
It was demonstrated in Levkov:2016rkk that the scale-invariant solution (14) with (18) is an attractor of the nonrelativistic system. This indicates that a sufficiently dense boson star is dragged towards a core collapse along the unstable branch.
@@
The second inequality in (22) is for a nonrelativistic evolution, and negligible sextic and higher self-interactions. These eventually break down as the collapse proceeds ($t\to 0^{-}$). For the case of an axion-like potential, the scale-invariant collapse is followed by a “bosenova” which emits relativistic axions and hence lowers the central density of the star Levkov:2016rkk. The diluted star then collapses again along the unstable branch, and repeats the whole process.
@@ 3.1 A toy model
Second, we have not studied in depth the basin of attraction for the collapsing solution, and how spherically non-symmetric configurations evolve. However since the collapse is initially governed by the quartic interaction, the results in Ref. Levkov:2016rkk which partially demonstrated that the spherically symmetric collapsing solution is an attractor for axion stars, also apply to our case. If the star becomes sufficiently spherically symmetric during the scale-invariant collapse, then it is reasonable to expect that the system remains nearly spherically symmetric also during the subsequent evolution.
@@ 4.1 Transition to our universe
Boson stars can form in virialized halos through a condensation Kolb:1993zz; Levkov:2018kau; Eggemeier:2019jsu; Chen:2021oot or by a gravothermal collapse Ralegankar:2024zjd; they can also form from the collapse of large-amplitude density fluctuations Gorghetto:2024vnp. In any cases, the formation of boson stars (or at least their host halos) typically takes place in a universe dominated by nonrelativistic matter. Hence one may imagine a scenario that leads to the star-induced phase transition as follows: The $\phi$ particles are initially excited around the false vacuum, which eventually become nonrelativistic and dominate the universe. During this early matter-dominated epoch, the $\phi$ particles form boson stars through one of the above mechanisms. The stars then grow by accreting the surrounding particles Chan:2022bkz; Dmitriev:2023ipv, some of which reach the unstable branch, collapse, and explode as bubbles of true vacuum.
@@
We have also assumed a matter-dominated epoch driven by the same field producing the boson stars. If the stars or at least their host halos can form during a different epoch, the previous discussion in principle does not apply. For instance, one can imagine another matter component that is not bounded by $\phi$’s potential barrier to dominate the universe and induce structure formation. Alternatively, halos might form in a non-matter-dominated universe with the aid of additional long-range forces Flores:2020drq.
@@ 5 Summary & Conclusion
We have briefly discussed the phenomenological implications. In particular, we demonstrated that a false vacuum that is metastable against quantum tunneling and hence may seem to have a lifetime exceeding the age of the universe, can actually be destabilized by a boson star collapse. At this point there are two distinct scenarios: (i) our world lives in the true vacuum, or (ii) our world lives in the false vacuum. Regarding the case (i), for the simplest toy scenario we showed that boson stars formed in a matter-dominated universe in the false vacuum can only collapse into subcritical bubbles, and hence do not trigger a phase transition. It would be interesting to see if a similar situation also arises for other boson potentials and star formation scenarios. For (ii), the false vacuum energy is normalized to the observed value of the cosmological constant and is very small. In this case, boson stars can form and trigger a destruction of the otherwise metastable universe. The requirement that such dangerous stars do not form can be used to constrain cosmological models with multiple vacua, such as those with multi-axion couplings or gravitational corrections generating superposed cosine potentials Graham:2015cka; Choi:2015fiu; Kaplan:2015fuy; Higaki:2016yqk; Kobayashi:2018nzh; DiLuzio:2021gos.
We also comment on gravitational waves emitted from the collision of bubbles. The frequency is set by the initial distance between the bubbles upon nucleation, which is roughly the Hubble radius in conventional first-order phase transition scenarios. However in our mechanism, the bubble separation should be much smaller since it corresponds to the interstellar distance. Hence for phase transitions happening at the same energy scale, we expect star-induced bubbles to produce gravitational waves with much higher frequencies than in conventional scenarios. On top of this, there may be additional gravitational wave emission from mergers of boson stars Croon:2018ftb; Chung-Jukko:2024hod, and from fluctuations of the expanding bubble wall Blum:2024hcs.
@@ Acknowledgements.
We thank Motoo Suzuki for initial collaboration. We also also thankful to Francescopaolo Lopez, Pranjal Ralegankar, Chen Sun, and Michael Zantedeschi for helpful discussions. This work was supported in part by the European Union - NextGenerationEU through the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W), and INFN initiatives TAsP and APINE. T.K. also acknowledges support from JSPS KAKENHI (JP22K03595).
@@ Appendix A O(3)-invariant bounce solution
In this appendix we discuss the critical radius of a true vacuum bubble in terms of an O(3)-invariant bounce solution. These static field configurations correspond to a saddle point of the energy functional, and under small perturbations can either collapse or expand. Our discussion is purely classical, however the computation is essentially the same as for critical bubble formation at finite temperatures due to thermal jumps Linde:1980tt.
@@
Interpreting the radial variable $r$ as time, then the equation (56) describes the motion of a particle moving along an inverted potential V-V, being also subject to a frictional force of $(2/r)(d\phi/dr)$. The boundary conditions further impose that the particle is initially ($r=0$) at rest, then subsequently starts to roll and reaches $\phi=0$ in the asymptotic future ($r\to\infty$). Here, since $\phi=0$ is a local maximum of the inverted potential, to reach it the particle needs to be initially located in the potential well around the true vacuum. Moreover, due to the presence of the frictional force, the initial position needs to fulfill $V(\phi_{B}(0))<V(0)$. We refer to $\phi_{B}(r)$ as the “bounce” solution. This can be numerically obtained for generic potentials using a shooting method, or by using existing codes such as FindBounce Guada:2020xnz and CosmoTransitions Wainwright:2011kj.
The bounce solution is unstable under deformations. For example we can look at the following one-parameter field configurations $\phi_{\kappa}(r)=\phi_{B}(\kappa r)$ with $\kappa>0$ (hence $\phi_{\kappa=1}=\phi_{B}$) Coleman:1985rnk. Noting that the bounce solution extremizes the Hamiltonian for $\phi$,

### v2

@@ Abstract
We present a new classical mechanism for nucleation of bubbles of true vacuum. The mechanism arises when dense boson stars form in the false vacuum. As the boson stars collapse due to attractive self-interactions, the field inside the star cores is enhanced beyond the potential barrier. Subsequently the stars explode as true vacuum bubbles, and induce a cosmological phase transition. The mechanism raises the possibility that a vacuum that is stable against quantum tunneling can be vulnerable to “astrophysical” processes.
@@ 1 Introduction
Recent advances in gravitational wave astronomy Barausse and others (2020); Afzal and others (2023) have sparked interest in the physics of cosmological first-order phase transitions Coleman (1977); Linde (1981), potentially leading to stochastic gravitational wave signals that are observable in current and future experiments. These experiments can reach frequencies up to a few kHz, which correspond to transition scales much higher than the physics scales reachable by collider experiments. [^1] A present-day frequency of kHz is tied to a transition scale of $\sim 10^{10}\,\mathrm{GeV}$ in conventional scenarios. This relation however is modified in the mechanism we will present. The signal details depend strongly on the physics behind the phase transition, which usually proceeds via quantum tunnelling Coleman (1977) or thermal fluctuations Linde (1981, 1980) creating supercritical bubbles which later expand and collide. In this paper, we would like to present a new classical mechanism for the creation of supercritical bubbles from the collapse of a boson star.
Boson stars Kaup (1968); Ruffini and Bonazzola (1969) are clumps of Bose-Einstein condensate that are held in equilibrium by gravity or attractive self-interactions. They are known to arise in various early universe scenarios Colpi et al. (1986); Liddle and Madsen (1992); Kolb and Tkachev (1993); Khlebnikov (2000); Sikivie and Yang (2009); Chavanis (2011); Chavanis and Delfini (2011); Chavanis (2025, 2018); Erken et al. (2012); Davidson and Elmer (2013); Davidson (2015); Guth et al. (2015); Eby et al. (2016); Levkov et al. (2017); Schiappacasse and Hertzberg (2018); Levkov et al. (2018); Eggemeier and Niemeyer (2019); Chen et al. (2022); Chan et al. (2024); Dmitriev et al. (2024); Gorghetto et al. (2024); Ralegankar et al. (2025), and their physics has been extensively discussed especially in the context of axions (see e.g. Marsh (2016); Zhang (2019) for reviews). One of the important consequences of boson stars is that their evolutionary track can end with a collapse, which drives the energy density in the core of the star to very large values. This is indeed the case for axions, whose attractive self-interaction renders axion stars unstable against collapse; the axion star collapse has been shown Levkov et al. (2017) to be accompanied by emission of relativistic axions from the high-density core.
@@
Generally speaking, the process we investigate can be viewed as a true vacuum bubble creation driven by field–interaction dynamics. In this context, we would like to mention Refs. Easther et al. (2009); Giblin et al. (2010), where a different type of field dynamics also leads to classical bubble production. In those works, large field excursions are generated by bubble collisions, which in turn result in the classical creation of new bubbles in minima lying below those of the parent bubbles. However, unlike the mechanism discussed in the present paper, the effect found in Easther et al. (2009); Giblin et al. (2010) operates only in models featuring a sequence of at least two false vacua arranged in a hierarchical structure. Finally, let us also mention other studies that are somewhat similar in spirit, where phase transitions are triggered by various types of nucleation sites Hosotani (1983); Preskill and Vilenkin (1993), such as neutron stars Balkin et al. (2022, 2023), black holes Hiscock (1987); Gregory et al. (2014); Mukaida and Yamada (2017); Shkerin and Sibiryakov (2021), domain walls Blasi and Mariotti (2022); Blasi et al. (2023); Agrawal et al. (2024), cosmic strings Yajnik (1986); Blasi and Mariotti (2025), and monopoles Steinhardt (1981); Kumar et al. (2010); Agrawal and Nee (2022).
The paper is structured as follows: in Section 2 we briefly review dense boson stars, focusing on their collapse according to the self-similar solution found by Levkov et al. (2017). In Section 3, we present the main result of this work and show for an explicit potential, how the boson star collapse can trigger the phase transition. In Section 4, we comment on phenomenological applications of this mechanism. Finally in Section 5 we discuss our findings and provide with our concluding remarks. We also give a general discussion about the critical radius of a true vacuum bubble in Appendix A.
@@ 2 Boson stars: a review
Let us start by reviewing the basic properties of boson stars, which are chunks of Bose-Einstein condensate bound by gravity or attractive self-interactions. Some of the discussions here follow those of Guth et al. (2015); Levkov et al. (2017); Schiappacasse and Hertzberg (2018).
@@ 2.1 Stable and unstable branches
where $r=\lvert\bm{x}\rvert$. Then the integrals in (7) can be performed to yield [^3] For exact star solutions of the GPP equations, the terms in the Hamiltonian (9) can be corrected by order-unity factors, whose values also depend on the detailed definitions of the star radius and mass Ruffini and Bonazzola (1969); Membrado et al. (1989); Visinelli et al. (2018). However the approximation (8) is good enough for order-of-magnitude estimates.
@@ 2.2 Boson star collapse
We now focus on star configurations in the vicinity of the unstable branch, where gravity can be ignored. Then the GPP equations (6) reduce to a nonlinear Schrödinger-type equation, which has been well studied in condensed matter systems Zakharov and Kuznetsov (2012),
@@
As noted in Zakharov and Kuznetsov (2012); Levkov et al. (2017), this equation is invariant under the scaling,
@@
where $A$ and $B$ are complex constants. The second term should vanish for the Hamiltonian to be finite. Numerically solving (15) under these boundary conditions fix the parameters as $w\simeq 0.54$ and $\lvert A\rvert\simeq 2.85$ Levkov et al. (2017). Moreover, the absolute value of $\chi$ thus obtained is well approximated by the following piece-wise function,
@@
Two comments are in order here. First, it is the star core whose mass-radius relation takes a similar form as that of the unstable branch. For the whole star, since it acquires a non-uniform density during the collapse, the discussion below (8) no longer applies. Moreover, considering the scale-invariant profile (19) to be cut off at the star radius $R_{\rm{star}}$ ($\gg R$), the total mass of the star is much larger than the core mass: $M_{\rm{star}}\simeq(3R_{\rm{star}}/R)M\gg M$. Second, it was demonstrated in Levkov et al. (2017) that the scale-invariant solution is an attractor. This indicates that a sufficiently dense boson star is dragged towards the scale-invariant core collapse.
@@
The second inequality in (22) is for a nonrelativistic evolution, and negligible sextic and higher self-interactions. These eventually break down as the collapse proceeds ($t\to 0^{-}$). For the case of an axion-like potential, the scale-invariant collapse is followed by a “bosenova” which emits relativistic axions and hence lowers the central density of the star Levkov et al. (2017). The diluted star then collapses again along the unstable branch, and repeats the whole process.
@@ 3.1 A toy model
Second, we have not studied in depth the basin of attraction for the collapsing solution, and how spherically non-symmetric configurations evolve. However since the collapse is initially governed by the quartic interaction, the results in Ref. Levkov et al. (2017) which partially demonstrated that the spherically symmetric collapsing solution is an attractor for axion stars, also apply to our case. If the star becomes sufficiently spherically symmetric during the scale-invariant collapse, then it is reasonable to expect that the system remains nearly spherically symmetric also during the subsequent evolution.
@@ 4.1 Transition to our universe
Boson stars can form in virialized halos through a condensation Kolb and Tkachev (1993); Levkov et al. (2018); Eggemeier and Niemeyer (2019); Chen et al. (2022) or by a gravothermal collapse Ralegankar et al. (2025); they can also form from the collapse of large-amplitude density fluctuations Gorghetto et al. (2024). In any cases, the formation of boson stars (or at least their host halos) typically takes place in a universe dominated by nonrelativistic matter. Hence one may imagine a scenario that leads to the star-induced phase transition as follows: The $\phi$ particles are initially excited around the false vacuum, which eventually become nonrelativistic and dominate the universe. During this early matter-dominated epoch, the $\phi$ particles form boson stars through one of the above mechanisms. The stars then grow by accreting the surrounding particles Chan et al. (2024); Dmitriev et al. (2024), some of which reach the unstable branch, collapse, and explode as bubbles of true vacuum.
@@
We have also assumed a matter-dominated epoch driven by the same field producing the boson stars. If the stars or at least their host halos can form during a different epoch, the previous discussion in principle does not apply. For instance, one can imagine another matter component that is not bounded by $\phi$’s potential barrier to dominate the universe and induce structure formation. Alternatively, halos might form in a non-matter-dominated universe with the aid of additional long-range forces Flores and Kusenko (2021).
@@ 5 Summary & Conclusion
We have briefly discussed the phenomenological implications. In particular, we demonstrated that a false vacuum that is metastable against quantum tunneling and hence may seem to have a lifetime exceeding the age of the universe, can actually be destabilized by a boson star collapse. At this point there are two distinct scenarios: (i) our world lives in the true vacuum, or (ii) our world lives in the false vacuum. Regarding the case (i), for the simplest toy scenario we showed that boson stars formed in a matter-dominated universe in the false vacuum can only collapse into subcritical bubbles, and hence do not trigger a phase transition. It would be interesting to see if a similar situation also arises for other boson potentials and star formation scenarios. For (ii), the false vacuum energy is normalized to the observed value of the cosmological constant and is very small. In this case, boson stars can form and trigger a destruction of the otherwise metastable universe. The requirement that such dangerous stars do not form can be used to constrain cosmological models with multiple vacua, such as those with multi-axion couplings or gravitational corrections generating superposed cosine potentials Graham et al. (2015); Choi and Im (2016); Kaplan and Rattazzi (2016); Higaki et al. (2016); Kobayashi and Ferreira (2018); Di Luzio et al. (2021).
We also comment on gravitational waves emitted from the collision of bubbles (see for a recent review Athron et al. (2024)). The frequency is set by the initial distance between the bubbles upon nucleation, which is roughly the Hubble radius in conventional first-order phase transition scenarios. However in our mechanism, the bubble separation should be much smaller since it corresponds to the interstellar distance. Hence for phase transitions happening at the same energy scale, we expect star-induced bubbles to produce gravitational waves with much higher frequencies than in conventional scenarios. On top of this, there may be additional gravitational wave emission from mergers of boson stars Croon et al. (2018); Chung-Jukko et al. (2024), and from fluctuations of the expanding bubble wall Blum and Mirbabayi (2024).
@@ Acknowledgements.
We thank Motoo Suzuki for initial collaboration. We are also thankful to Francescopaolo Lopez, Pranjal Ralegankar, Chen Sun, and Michael Zantedeschi for helpful discussions. This work was supported in part by the European Union - NextGenerationEU through the PRIN Project “Charting unexplored avenues in Dark Matter” (20224JR28W), and INFN initiatives TAsP and APINE. T.K. also acknowledges support from JSPS KAKENHI (JP22K03595).
@@ Appendix A O(3)-invariant bounce solution
In this appendix we discuss the critical radius of a true vacuum bubble in terms of an O(3)-invariant bounce solution. These static field configurations correspond to a saddle point of the energy functional, and under small perturbations can either collapse or expand. Our discussion is purely classical, however the computation is essentially the same as for critical bubble formation at finite temperatures due to thermal jumps Linde (1981).
@@
Interpreting the radial variable $r$ as time, then the equation (56) describes the motion of a particle moving along an inverted potential V-V, being also subject to a frictional force of $(2/r)(d\phi/dr)$. The boundary conditions further impose that the particle is initially ($r=0$) at rest, then subsequently starts to roll and reaches $\phi=0$ in the asymptotic future ($r\to\infty$). Here, since $\phi=0$ is a local maximum of the inverted potential, to reach it the particle needs to be initially located in the potential well around the true vacuum. Moreover, due to the presence of the frictional force, the initial position needs to fulfill $V(\phi_{B}(0))<V(0)$. We refer to $\phi_{B}(r)$ as the “bounce” solution. This can be numerically obtained for generic potentials using a shooting method, or by using existing codes such as FindBounce Guada et al. (2020) and CosmoTransitions Wainwright (2012).
The bounce solution is unstable under deformations. For example we can look at the following one-parameter field configurations $\phi_{\kappa}(r)=\phi_{B}(\kappa r)$ with $\kappa>0$ (hence $\phi_{\kappa=1}=\phi_{B}$) Coleman (1985). Noting that the bounce solution extremizes the Hamiltonian for $\phi$,